# Generating Hypotheses of Dynamic Causal Graphs in Neuroscience: Leveraging Generative Factor Models of Observed Time Series

### v1

@@ C.6 Ablation Analyses
In Table 8 we provide the details of our ablation experiments summarized in Table 3 of the main paper. With the exception of our $\rho=0$ ablation experiments, we see that ablation scores are worse - in many cases by a large margin - than those of the full REDCLIFF-S model, suggesting that these hyperparameters play an important role in REDCLIFF-S’ performance.
The $\rho=0$ ablation experiments tell a more nuanced story, with performance degrading marginally for three of the five systems tested and actually increasing when $\rho$ was ablated for two of the systems. These observations are somewhat unsurprising for a few reasons. Firstly, we reiterate that we only tuned the hyperparameters - including $\rho$ on a single synthetic system configuration due to time constraints, so it is unsurprising that at least one of our parameter settings is sub-optimal in certain cases. Secondly, we note that there appears to be a pattern in Table 8 where the systems with sparser edges and fewer factors (esp. the 6-2-2 and 12-11-2 Synthetic Systems) tend to perform worse after setting $\rho=0$, whereas those with more edges and factors (esp. the 6-4-2 and 12-11-5 Synthetic Systems) see improvements when $\rho=0$. This makes intuitive sense, since state graphs in systems with more edges and factors are simply more likely to share similar causal relationships than those with sparser connections and fewer states. Finally, the CosSim loss term which $\rho$ modulates was introduced into REDCLIFF’s training process after we observed that doing so correlated loosely with improved ROC-AUC performance at test time in preliminary experiments (see Appendix B.3); since the ROC-AUC and F1 metrics are fundamentally different, the correlated relationship between minimized CosSim and ROC-AUC may not always indicate improved F1 performance. In summary, it would seem that $\rho$ is more sensitive than other hyperparameters to properties of the system being modeled and should thus be carefully tuned to suit the particular target system (and possibly dropped entirely for sufficiently dense or expressive systems).
@@ Appendix D Datasets Used: Additional Details
In this section we report various hyperparameters used in curating/preparing the datasets used in our experiments. Supplemental Table 9 presents hyperparameters for the D4IC dataset (see Section 4.3). Supplemental Table 10 presents hyperparameters for the Synthetic Systems Dataset (see Section 4.2). Supplemental Table 11 presents hyperparameters for the Region-Averaged TST 100 Hz dataset (see Section 5). We briefly discuss the Region-Averaged SP 100 Hz dataset in Appendix D.2.
@@
Table 10: Synthetic Systems Dataset (Sections 4.1-4.2) - hyperparameters across experiments. Of particular note, repeat number 3 (indexed from 0) of the $\mathrm{n_{c}}=12$-node, $\mathrm{n_{e}}=33$-edge (inter-variable), $\mathrm{n_{k}}=3$-factor synthetic system dataset was used for grid-searching hyperparameters across our Synthetic Systems experiments. Generally speaking, we find that the greater the number of edges per node in the system, the more ‘complex’ the Granger causal estimation task.
@@ D.1 Preprocessing the TST 100 Hz Datasets
As mentioned in the main paper, we applied some preprocessing steps to the publicly available TST dataset (Carlson et al., 2023) in order to prepare it for our experiments. Specifically, we first sorted the publicly available local field potential (LFP) recordings and associated behavioral label files according to the identifier of the mouse being observed; this was done in order to ensure all mice were equally represented in each data repeat and so that different mice could be designated as ‘hold-out’ subjects for each repeat. We then iterated through each subject mouse’s files and performed the following operations on each recording:
@@
  • Marked any time point for which the LFP value was more than 15 median absolute deviations above the median value of the (filtered - see below) recorded signal as outliers (i.e. ‘nan’ values);
@@
Once a recording file had been run through the above steps, we drew samples evenly across the portions of the recording representing the a) Home Cage (HC), b) Open Field (OP), and c) Tail Suspended (TS) behavioral paradigms. Each sample was drawn such that it was guaranteed to not have any marked-outlier time points. Once a sample had been drawn, a one-hot vector label corresponding to the behavioral paradigm was attached to it and the sampled signal was down-sampled to reduce computational overhead in later experiments.
@@ Appendix E Model Hyperparameters Used in Experiments: Additional Details
In this section we report hyperparameters used to configure and train the models used in our experiments. Hyperparameters for the cLSTM models which appear in this paper (see Sections 4.2 and 4.3) are given in Supplemental Table 12. Hyperparameters for the cMLP models which appear in this paper (see Sections 4.2 and 4.3) are given in Supplemental Table 13. Hyperparameters for the DGCNN models which appear in this paper (see Sections 4.2 and 4.3) are given in Supplemental Table 14. Hyperparameters for the DYNOTEARS models which appear in this paper (see Section 4.3) are given in Supplemental Table 15. Hyperparameters for the NAVAR-R (cLSTM) models which appear in this paper (see Section 4.3) are given in Supplemental Table 16. Hyperparameters for the NAVAR-P (cMLP) models which appear in this paper (see Section 4.3) are given in Supplemental Table 17. Hyperparameters for the dCSFA models which appear in this paper (see Sections 4.2 and 4.3) are given in Supplemental Table 18. Hyperparameters for the REDCLIFF-S (cMLP) models which appear in this paper (see Sections 4.2, 4.3, and 5) are given in Supplemental Table 19 and 20.
@@
Table 20: REDCLIFF-S - hyperparameters across experiments on real-world data. Note that for the Region-Averaged TST 100 Hz dataset we had $n_{\mathrm{c}}=12$ while for the Regaion Averaged SP dataset we had $n_{\mathrm{c}}=9$.
@@ E.2 Hyperparameters Represented by Functions
We noticed in preliminary experiments that some hyperparameters were highly sensitive to certain aspects of the dynamical systems being modeled; furthermore, we noticed these sensitivities could be described by fairly simple functions of system aspects. Rather than set a fixed value for these model hyperparameters, we tuned ‘meta-hyperparameters’ (e.g. coefficients) of the functions for these sensitive hyperparameters to a particular dataset, and then re-computed the output of their functions each time the ‘tuned’ model was transferred to a new target dynamical system. We report the tuned functional form (with tuned meta-hyperparameters) in place of any fixed value for this type of sensitive hyperparameter.

### v2

@@ C.6 Ablation Analyses
In Table 8 we provide details of our ablation experiments summarized in Table 3 of the main paper. With the exception of our $\rho=0$ ablation experiments, we see that the ablation scores are worse - in many cases by a large margin - than those of the full REDCLIFF-S model, suggesting that these hyperparameters play an important role in the performance of REDCLIFF-S.
The $\rho=0$ ablation experiments tell a more nuanced story, with performance degrading marginally for three of the five systems tested and actually increasing when $\rho$ was ablated for two of the systems. These observations are somewhat unsurprising for a few reasons. Firstly, we reiterate that we only tuned the hyperparameters - including $\rho$ on a single synthetic system configuration due to time constraints, so it is unsurprising that at least one of our parameter settings is suboptimal in certain cases. Secondly, we note that there appears to be a pattern in Table 8 where the systems with relatively fewer edges and factors (esp. the 6-2-2 and 12-11-2 Synthetic Systems) tend to perform worse after setting $\rho=0$, whereas those with more edges and factors (esp. the 6-4-2 and 12-11-5 Synthetic Systems) see improvements when $\rho=0$. This makes intuitive sense, since state graphs in systems with more edges and factors are simply more likely to share similar causal relationships than those with fewer connections and states. Finally, the CosSim loss term modulated by $\rho$ was introduced into the REDCLIFF training process after we observed that doing so was loosely correlated with improved ROC-AUC performance at test time in preliminary experiments (see Appendix B.3); since the ROC-AUC and F1 metrics are fundamentally different, the correlated relationship between minimized CosSim and ROC-AUC may not always indicate improved F1 performance. In summary, it would seem that $\rho$ is more sensitive than other hyperparameters to properties of the system being modeled and should thus be carefully tuned to suit the particular target system (and possibly dropped entirely for sufficiently dense or expressive systems).
@@ Appendix D Datasets Used: Additional Details
In this section, we report various hyperparameters used in curating/preparing the datasets used in our experiments. Supplemental Table 9 presents hyperparameters for the D4IC dataset (see Section 4.3). Supplemental Table 10 presents hyperparameters for the Synthetic Systems Dataset (see Section 4.2). Supplemental Table 11 presents hyperparameters for the Region-Averaged TST 100 Hz dataset (see Section 5). We briefly discuss the Region-Averaged SP 100 Hz dataset in Appendix D.2.
@@
Table 10: Synthetic Systems Dataset (Sections 4.1-4.2) - hyperparameters across experiments. Of particular note, repeat number 3 (indexed from 0) of the $\mathrm{n_{c}}=12$-node, $\mathrm{n_{e}}=33$-edge (inter-variable), $\mathrm{n_{k}}=3$-factor synthetic system dataset was used to perform grid searches over hyperparameters across our Synthetic Systems experiments. Generally speaking, we find that the greater the number of edges per node in the system, the more ‘complex’ the Granger causal estimation task.
@@ D.1 Preprocessing the TST 100 Hz Datasets
As mentioned in the main paper, we applied some preprocessing steps to the publicly available TST dataset (Carlson et al., 2023) to prepare it for our experiments. Specifically, we first sorted the publicly available local field potential (LFP) recordings and associated behavioral label files according to the identifier of the mouse being observed; this was done to ensure all mice were equally represented in each data repeat and so that different mice could be designated as ‘holdout’ subjects for each repeat. We then iterated through each subject mouse’s files and performed the following operations on each recording:
@@
  • Marked any time point for which the LFP value was more than 15 median absolute deviations above the median value of the (filtered - see below) recorded signal as outliers (i.e., ‘nan’ values);
@@
Once a recording file had been run through the above steps, we drew samples evenly across the parts of the recording representing the a) Home Cage (HC), b) Open Field (OP), and c) Tail Suspended (TS) behavioral paradigms. Each sample was drawn in such a way that it was guaranteed to have no marked outlier time points. Once a sample had been drawn, a one-hot vector label corresponding to the behavioral paradigm was attached to it, and the sampled signal was downsampled to reduce computational overhead in later experiments.
@@ Appendix E Model Hyperparameters Used in Experiments: Additional Details
In this section, we report hyperparameters used to configure and train the models used in our experiments. Hyperparameters for the cLSTM models that appear in this paper (see Sections 4.2 and 4.3) are given in Supplemental Table 12. Hyperparameters for the cMLP models that appear in this paper (see Sections 4.2 and 4.3) are given in Supplemental Table 13. Hyperparameters for the DGCNN models that appear in this paper (see Sections 4.2 and 4.3) are given in Supplemental Table 14. Hyperparameters for the DYNOTEARS models that appear in this paper (see Section 4.3) are given in Supplemental Table 15. Hyperparameters for the NAVAR-R (cLSTM) models that appear in this paper (see Section 4.3) are given in Supplemental Table 16. Hyperparameters for the NAVAR-P (cMLP) models that appear in this paper (see Section 4.3) are given in Supplemental Table 17. Hyperparameters for the dCSFA models that appear in this paper (see Sections 4.2 and 4.3) are given in Supplemental Table 18. Hyperparameters for the REDCLIFF-S (cMLP) models that appear in this paper (see Sections 4.2, 4.3, and 5) are given in Supplemental Tables 19 and 20.
@@
Table 20: REDCLIFF-S - hyperparameters across experiments on real-world data. Note that for the Region-Averaged TST 100 Hz dataset we had $n_{\mathrm{c}}=12$ while for the Region Averaged SP dataset we had $n_{\mathrm{c}}=9$.
@@ E.2 Hyperparameters Represented by Functions
We noticed in preliminary experiments that some hyperparameters were highly sensitive to certain aspects of the dynamical systems being modeled; furthermore, we noticed these sensitivities could be described by fairly simple functions of system aspects. Rather than set a fixed value for these model hyperparameters, we tuned ‘meta-hyperparameters’ (e.g., coefficients) of the functions for these sensitive hyperparameters to a particular dataset, and then recomputed the output of their functions each time the ‘tuned’ model was transferred to a new target dynamical system. We report the tuned functional form (with tuned meta-hyperparameters) in place of any fixed value for this type of sensitive hyperparameter.